# score-based generative model SNR signal-to-noise ratio

### v2

@@ V-B Dereverberation
Zero-phase - 1.31 0.68 -34.6 1.73 1.86 1.42 14.9% 1.19 - -
SFM Euler1 1 1.58 0.58 1.7 1.90 1.70 2.05 61.8% 6.72 32 27.9M
SFM Euler5 5 4.24 0.97 -1.7 4.3 3.26 4.13 3.7% 0.76 32 27.9M
SFM Midpoint2 4 4.05 0.96 -2.5 4.3 3.15 4.15 3.7% 0.68 32 27.9M
SFM LRK5 5 4.22 0.97 -2.3 4.4 3.27 4.11 3.0% 0.61 32 27.9M
RTISI [beauregard2005efficient] 50 3.08 0.90 -28.3 3.28 2.76 3.03 5.2% 0.71 32 0
RTISI-DM [peer2024flexible] 50 3.35 0.91 -27.0 3.86 2.83 3.56 4.9% 0.73 32 1
FM Euler5 5 4.38 0.98 -1.2 4.37 3.25 4.10 2.9% 0.65 $\infty$ 38.7M
TABLE VIII: Mel vocoding task on EARS-WHAM v2 test set clean utterances. HiFi-GAN refers to the 16 kHz model trained on LibriTTS data [zen2019libritts], available in SpeechBrain [ravanelli2024speechbrain]. For RTISI-DM, NFE means the number of algorithm iterations. Best within a group in bold, second best underlined, worse than input in red. We refer the reader to our prior work [welker2025realtime] for a more in-depth evaluation on this task.
@@
$M^{\dagger}$ + Zero-phase - 1.28 0.63 -38.9 1.43 1.07 1.21 35.1% 2.22 -
SFM Euler1 1 1.35 0.36 -5.6 1.18 1.23 1.35 86.7% 7.82 32 27.9M
SFM Euler5 5 4.10 0.96 -10.1 4.31 3.11 4.14 4.7% 1.01 32 27.9M
SFM Midpoint2 4 3.92 0.94 -10.2 4.28 2.99 4.17 4.5% 0.89 32 27.9M
SFM LRK5 5 4.14 0.96 -10.4 4.34 3.15 4.15 3.6% 0.80 32 27.9M
$M^{\dagger}$ + RTISI-DM [peer2024flexible] 50 2.97 0.88 -29.7 2.51 1.92 2.58 5.8% 0.82 32 1
FM Euler5 5 4.34 0.97 -9.9 4.35 3.05 4.15 4.3% 1.04 $\infty$ 38.7M
HiFi-GAN [kong2020hifigan] 1 2.99 0.90 -29.9 4.21 3.02 3.91 5.4% 0.77 236 13.9M
For dereverberation, the results are shown in Table IV. SFM achieves a consistent improvement over the reverberant audio, with Euler5 generally performing best and the learned RK5 solver not yielding a clear improvement. The non-streaming FM baseline shows a clear advantage, particularly in PESQ and WER. We argue that this is expectable, since future information is especially useful to estimate RIR characteristics and suppress reverberation.
V-C Codec post-filtering
For codec post-filtering, we list the results in Table V. SFM can greatly improve both intrusive and non-intrusive metrics over the plain Lyra V2 decoder, at the cost of a small increase in WER. The learned Runge-Kutta scheme fails to yield an improvement here except in LSD. We conjecture that (19) is suboptimal here since the Lyra decoder outputs already have good phoneme and high-frequency preservation, but we leave other loss choices to future work. Notably, the non-streaming FM baseline shows only marginal quality improvements over SFM at five Euler steps here, which may be related to the streaming nature of the Lyra V2 codec itself.
V-D Bandwidth extension
We list the bandwidth extension metrics in Table VI. The Euler1 solver is now clearly the worst option, with a significant gap in non-intrusive metrics and WER compared to higher-NFE solvers. Notably, the learned RK5 solver performs best, with a clear improvement in all metrics except PESQ and SI-SDR. LRK5 also clearly improves WER, suggesting that it reconstructs the semantic content better. This confirms the usefulness of our proposal in this highly ambiguous restoration problem. PESQ may be misleading here, as the bandlimited audios receive the best PESQ scores. The improvements from the learned solver are also supported by the listening experiment Fig. 2(b), where it receives the highest scores.
V-E STFT phase retrieval / Mel vocoding
For STFT PR and Mel vocoding, see Tables VII and VIII, which are both highly ill-posed non-linear inverse problems, the behaviors of the solvers are similar as for BWE, but the differences are more pronounced. Euler1 produces unusable estimates here, with worse metrics even than naive zero-phase estimates, and the learned solvers show a clear advantage, particularly in WER and LSD.
V-F Computations and runtimes
TABLE IX: FLOP and RTF measurements on an RTX 4080 Laptop GPU. NN is the number of DNN calls per frame. FLOPs are per-frame multiplied with the number of frames per second, streaming RTF is relative to each model’s per-frame runtime budget (frame shift), and offline RTF is determined for 1-second inputs.
@@ V-F Computations and runtimes
Diffusion Buffer [lay2025diffusionjpre] 8810.4 0.646 0.007
DEMUCS [defossez2020realtime] 10.2 0.221 0.004
HiFi-Stream [dmitrieva2025hifi] 6.7 0.098 0.012
DeepFilterNet3 [schroeter2023deepfilternet3] 0.3 0.321 0.008
@@
In Table IX we list the determined GFLOPs per second, the streaming RTF defined as $\frac{t_{\mathrm{proc,fr}}}{H_{\text{syn}}/f_{s}}$ for each single frame where $f_{s}$ is the sampling frequency and $t_{\mathrm{proc,fr}}$ is the processing time per frame [defossez2020realtime], and offline RTF $\frac{t_{\mathrm{proc}}}{1\,\text{sec}}$ where $t_{\mathrm{proc}}$ is the time taken to process one 1-second utterance in a single model call. We can see that offline RTF consistently and severely underestimates the streaming RTF, and that FLOPs have no simple relation to streaming RTF, e.g., the Diffusion Buffer [lay2025diffusionjpre] attains smaller streaming and offline RTFs than Stream.FM for 5 NFE but spends substantially more FLOPs.
V-G Model weight compression for Mel vocoding
@@ VI Conclusion
In this work, we have presented Stream.FM, a streaming generative method for general speech restoration tasks which can run in real-time on a consumer GPU. We detailed our architecture, inference scheme, and other optimizations needed to achieve this real-time capability, showed state-of-the-art performance for generative streaming methods through metrics and listening experiments, and proposed novel ideas to optimize quality or compute in a low-NFE setting. We hope that our contributions and public codebase can help to close the gap between generative and predictive speech restoration models in real-time settings, and support the research community towards further developments in this field.
@@ VII Acknowledgements
[Uncaptioned image] Simon Welker received a B.Sc. in Computing in Science (2019) and M.Sc. in Bioinformatics (2021) from University of Hamburg, Germany. He is currently a PhD student in the labs of Prof. Timo Gerkmann (Signal Processing, University of Hamburg) and Prof. Henry N. Chapman (Center for Free-Electron Laser Science, DESY, Hamburg), researching machine learning techniques for solving inverse problems that arise in speech processing and X-ray imaging. He received the VDE ITG award 2024.
[Uncaptioned image] Bunlong Lay obtained a B.Sc. and M.Sc. in Mathematics in 2015 and 2017 from the University of Bonn, Germany. He subsequently joined the research institute Fraunhofer FKIE in Wachtberg Germany from 2018 until 2021, where he focused on research in the field of radar signal processing. In 2021 he started his Ph.D. at the University of Hamburg. Currently researching Diffusion-based models for Speech Enhancement for real-time applications. He received the VDE ITG award 2024.
[Uncaptioned image] Maris Hillemann obtained a B.Sc. in Computer Science in 2024 from the University of Hamburg, Germany. He is currently a master’s student in Computer Science at the University of Hamburg and a student assistant in the Signal Processing group of Prof. Timo Gerkmann.
[Uncaptioned image] Tal Peer received the B.Sc. degree in General Engineering Science (2016) and the M.Sc. degree in Electrical Engineering (2019) from the Hamburg University of Technology. He is currently pursuing a PhD with the Signal Processing group at the University of Hamburg. His research interests include phase-aware speech enhancement and phase retrieval for speech and audio applications.
[Uncaptioned image] Timo Gerkmann (S’08–M’10–SM’15) is a professor for Signal Processing at the Universität Hamburg, Germany. He has previously held positions at Technicolor Research & Innovation in Germany, the University of Oldenburg in Germany, KTH Royal Institute of Technology in Sweden, Ruhr-Universität Bochum in Germany, and Siemens Corporate Research in Princeton, NJ, USA. His main research interests are on statistical signal processing and machine learning for speech and audio applied to communication devices, hearing instruments, audio-visual media, and human-machine interfaces. Timo Gerkmann served as member of the IEEE Signal Processing Society Technical Committee on Audio and Acoustic Signal Processing and is currently a Senior Area Editor of the IEEE/ACM Transactions on Audio, Speech and Language Processing. He received the VDE ITG award 2022.
VIII Supplementary Material
In the following, we list additional technical details for our Stream.FM models, which are not necessary for the method itself but are useful to implement real-time capable streaming inference.
VIII-A Functional frame-wise inference API
To implement our frame-wise inference scheme, we implement our model using PyTorch and PyTorch Lightning. we design an application programming interface (API) similar to the one proposed for CINs [hedegaard2022cins]. The official CINs implementation stores tracked states on each layer, making every layer stateful. This is problematic for our purposes, since we want to use only a single DNN instance $v_{\theta}$ with shared weights, but multiple state collections (one for each diffusion solver step).
@@ VIII-A Functional frame-wise inference API
VIII-B Minimizing overhead
In initial real-time experiments with our models for real-time streaming inference, we found that the processing time per frame quickly exceeded our 16 ms budget even for low NFE values such as 3. After careful inspection of profiler traces, we found that the processing was dominated by CPU-GPU overhead. To reduce this overhead as much as possible, we first tried using model compilation via torch.compile to merge CUDA kernels, which reduced the overhead to some extent but not enough to allow real-time inference. We then modified our code to use CUDA Graphs as implemented in PyTorch. By capturing the entire solver computation including the sequence of $N$ DNN calls in a single CUDA Graph and then replaying this CUDA Graph for every frame, we are able to realize up to 5 DNN calls per frame within the 16 ms time budget.
VIII-C Learned coefficients for custom learned low-NFE ODE solvers
In this section, we list the parameters $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ of the learned Runge-Kutta ODE solver schemes, see Section III.E in the main paper, for each of the six tasks investigated in the main paper.
VIII-C1 Speech enhancement
@@ VIII-C2 Dereverberation
VIII-C3 Codec post-filtering
@@ VIII-C4 Bandwidth extension
VIII-C5 STFT phase retrieval

### v3

@@ V-B Dereverberation
V-C Codec post-filtering
For codec post-filtering, we list the results in Table IV(b). SFM can greatly improve both intrusive and non-intrusive metrics over the plain Lyra V2 decoder, at the cost of a small increase in WER. The learned Runge-Kutta scheme fails to yield an improvement here except in LSD. We conjecture that the loss (18) is suboptimal here since the Lyra decoder outputs already have good phoneme and frequency envelope preservation. Using the PESQ-based loss (19) instead improves intrusive metrics (PESQ, ESTOI, SI-SDR) and WVMOS slightly over Euler5, but clearly degrades all other metrics. We suspect that this task is especially difficult to optimize further due to the complicated ambiguities induced by the heavy compression of the Lyra V2 codec. Notably, the non-streaming FM baseline shows only marginal quality improvements over SFM at five Euler steps here, which may be related to the streaming nature of the Lyra V2 codec itself. AnyEnhance [68], which was also trained for codec artifact removal but only for MP3-encoded speech, degrades all intrusive metrics and WER below the decoded audio, but improves all non-intrusive metrics and, from informal listening, does somewhat improve the perceived audio quality.
V-D Bandwidth extension
We list the bandwidth extension metrics in Table IV(c). The Euler1 solver is now clearly the worst option, with a significant gap in non-intrusive metrics and WER compared to higher-NFE solvers. The AnyEnhance baseline [68] attains very good non-intrusive metric results, but clearly falls short of both our SFM and our FM model in all intrusive metrics (PESQ, ESTOI, SI-SDR, LSD) and WER. Its SI-SDR value of 6.9 is particularly low, which may be related to AnyEnhance performing enhancement in a compressed token space. Notably, the learned RK5 solver performs best, with a clear improvement in all metrics except PESQ and SI-SDR. LRK5 based on (18) also clearly improves WER, suggesting that it reconstructs the semantic content better. This confirms the usefulness of our proposal in this highly ambiguous restoration problem. The improvements from this learned solver are also supported by the listening experiment Fig. 2(b), where it receives the highest scores. Note that PESQ may be misleading here, as the bandlimited audios receive the best PESQ scores, except for LRK5 with the PESQ-based loss (19). While this LRK5 variant achieves the best PESQ (3.71), it is worse than Euler5 in all non-intrusive metrics and LSD. This is reflected in the outputs of this solver lacking high-frequency detail, see our audio example website. Thus, the solver learned with the PESQ-based objective does not adequately solve the desired bandwidth extension task.
V-E STFT phase retrieval / Mel vocoding
For STFT PR and Mel vocoding, see Tables V(a) and V(b), which are both highly ill-posed non-linear inverse problems, the behaviors of the solvers are similar as for BWE, but the differences are more pronounced. Euler1 produces unusable estimates here, with worse metrics even than naive zero-phase estimates. The learned solvers show a clear advantage, particularly in WER and LSD. On STFT phase retrieval, both Stream.FM and the non-causal FM baseline perform better than the non-causal DiffPhase [40], except in WER and LSD, while using fewer parameters and 6 times fewer NFE. On Mel vocoding, our streaming models outperform HiFi-GAN on all metrics except LSD, see also our conference publication [61]. Overall, Stream.FM shows excellent performance in these two tasks, with PESQ (>4.1>4.1) and ESTOI ($\geq 0.96$) values approaching the optimum.
TABLE V: Metric evaluation for STFT phase retrieval and Mel vocoding on EWv2 test set clean utterances. RTISI-DM refers to [41] using the Difference Map algorithm [11] with hyperparameter $\beta=1.75$, which we grid-searched as the optimum for 50 iterations. RTISI has no parameters. DiffPhase [40] uses matched training, see Section IV-A5. For RTISI and RTISI-DM, NFE refers to the number of algorithm iterations. Best within a group bold, second best underlined, worse than input in red. $\ell_{\text{alg}}$ and $\ell_{\text{tot}}$ in milliseconds.
(a) STFT phase retrieval task with 50% overlap, and no lookahead for all methods except non-causal FM.
@@ V-E STFT phase retrieval / Mel vocoding
Zero-phase - 1.31 0.68 -34.6 1.73 1.86 1.42 14.9% 1.19 - - -
RTISI [3] 50 3.08 0.90 -28.3 3.28 2.76 3.03 5.2% 0.71 32 48 0
RTISI-DM [41] 50 3.35 0.91 -27.0 3.86 2.83 3.56 4.9% 0.73 32 48 1
SFM Euler1 1 1.58 0.58 1.7 1.90 1.70 2.05 61.8% 6.72 32 48 27.9M
SFM Euler5 5 4.24 0.97 -1.7 4.30 3.26 4.13 3.7% 0.76 32 48 27.9M
SFM Midpoint2 4 4.05 0.96 -2.5 4.30 3.15 4.15 3.7% 0.68 32 48 27.9M
SFM LRK5 (18) 5 4.22 0.97 -2.3 4.40 3.27 4.11 3.0% 0.61 32 48 27.9M
FM Euler5 5 4.38 0.98 -1.2 4.37 3.25 4.10 2.9% 0.65 $\infty$ $\infty$ 38.7M
DiffPhase [40] 30 4.04 0.97 -25.3 4.29 3.19 3.96 2.6% 0.53 $\infty$ $\infty$ 65.6M
(b) Mel vocoding task. HiFi-GAN refers to the 16 kHz model trained on LibriTTS data [67], available in SpeechBrain [45]. We refer the reader to our prior work [61] for a more in-depth evaluation on this task.
[…]
$M^{\dagger}$ + Zero-phase - 1.28 0.63 -38.9 1.43 1.07 1.21 35.1% 2.22 - - -
$M^{\dagger}$ + RTISI-DM [41] 50 2.97 0.88 -29.7 2.51 1.92 2.58 5.8% 0.82 32 48 1
SFM Euler1 1 1.35 0.36 -5.6 1.18 1.23 1.35 86.7% 7.82 32 48 27.9M
SFM Euler5 5 4.10 0.96 -10.1 4.31 3.11 4.14 4.7% 1.01 32 48 27.9M
SFM Midpoint2 4 3.92 0.94 -10.2 4.28 2.99 4.17 4.5% 0.89 32 48 27.9M
SFM LRK5 (18) 5 4.14 0.96 -10.4 4.34 3.15 4.15 3.6% 0.80 32 48 27.9M
FM Euler5 5 4.34 0.97 -9.9 4.35 3.05 4.15 4.3% 1.04 $\infty$ $\infty$ 38.7M
HiFi-GAN [22] 1 2.99 0.90 -29.9 4.21 3.02 3.91 5.4% 0.77 236 252 13.9M
V-F Computations, runtimes, and memory
TABLE VI: FLOP and RTF measurements on an RTX 4080 Laptop GPU. NN is the number of DNN calls per frame. FLOPs are per-frame, multiplied with the number of frames per second, streaming RTF is relative to each model’s per-frame runtime budget (frame shift), and offline RTF is determined for 1-second inputs. Model variants as in Table II.
@@ V-F Computations, runtimes, and memory
Diffusion Buffer [28] 3529.8 0.475 0.007
Causal SEMamba [7] 42.5 0.867 0.016
DeepFilterNet3 [53] 0.3 0.321 0.008
DEMUCS [9] 10.2 0.221 0.004
CleanUMamba [12] n/a 0.278 0.004
HiFi-Stream [10] 6.7 0.098 0.012
@@
In Table VI we list, for our model and various streaming-capable baselines: (1) the determined GFLOPs per second, (2) the streaming RTF, defined as $\frac{t_{\mathrm{proc,fr}}}{H_{\text{syn}}/f_{s}}$ for each single frame where $f_{s}$ is the sampling frequency and $t_{\mathrm{proc,fr}}$ is the processing time per frame [9], and (3) the offline RTF $\frac{t_{\mathrm{proc}}}{1\,\text{sec}}$, where $t_{\mathrm{proc}}$ is the time taken to process one 1-second utterance in a single model call. We can see that offline RTF consistently and severely underestimates the streaming RTF, and that FLOPs have no simple relation to streaming RTF, e.g., the Diffusion Buffer [28] attains smaller streaming and offline RTFs than Stream.FM for 5 NFE but spends substantially more FLOPs. For CleanUMamba [12], no FLOPs measurement is available due to an incompatibility with the utilized FLOP counting method, see Section IV-D5, and we instead refer the reader to the GMACs values reported in the paper [12].
On a modern CPU (AMD Ryzen 7 9800X3D), we measure the streaming RTF of Stream.FM as $N\times 0.883$, which still permits real-time streaming with $N=1$ DNN evaluation and no initial predictor DNN, but precludes the use of any more than one model call, motivating the future search for more efficient DNN architectures.
Note that multiplying each model’s streaming RTF by its frame shift yields the effective processing latency $\ell_{\text{proc}}$, and adding this to the algorithmic latency $\ell_{\text{alg}}$ provides a lower bound for each model’s total latency $\ell_{\text{tot}}$, on this concrete hardware. However, as long as the streaming RTF is below 1, simply using $\ell_{\text{tot}}=W_{\text{syn}}+H_{\text{syn}}$ is closer to typical frame-by-frame implementations, see Section II-B.
Regarding memory usage in a real-time streaming setting, we analyze a streaming implementation of our SE model. We measure the total used GPU memory as $583+(45\cdot N)$ MiB. We further measure the memory used only by the tensors caching past activations, see Section III-B, as 15 MiB per solver step (plus 15 MiB for the initial predictor). These values permit deployment on any modern consumer PC, and potentially also some embedded devices.
V-G Learned Runge-Kutta solvers
Figure 3: Model error and curvature of the learned models in each task. Left: Normalized model error compared to the constant ground-truth velocity $u$ at each $\tau$, and learned Runge-Kutta solver time-points $\mathbf{c}$ learned with (18) for each task. Right: geometric curvature $\kappa$ [18] accumulated up until each indicated $\tau$. A fully straight velocity field would have zero curvature $\kappa$ even under (constant) model error.
TABLE VII: Changes in metrics for bandwidth extension by using an LRK solver trained with loss (18) instead of the Euler solver, compared at equal NFE. Improved values by LRK in green, worsened values in red, biggest improvement / smallest degradation in bold.
[…]
5 -0.35 -1.20 0.12 0.06 0.23 -1.8% -0.26
4 -0.34 -0.81 0.23 0.11 0.30 -1.8% -0.32
3 -0.47 -1.20 0.35 0.16 0.44 -2.3% -0.41
2 -0.20 -0.20 0.25 0.11 0.30 -0.2% -0.30
In our previous evaluations, we have seen that the LRK solvers can boost output quality both in objective metrics and in a listening test, but to which degree depends on the task and the loss used to train the solver. The LRK solvers are especially effective for BWE, STFT phase retrieval, and Mel vocoding, but do not yield such strong improvements for SE, dereverberation, and Lyra codec artifact removal. As a possible explanation for this inconsistent improvement, we propose that it is related to the curvature of the learned velocity field $v_{\theta}$ along the flow time $\tau$. To support this, we show an empirical estimation of the model error and the geometric curvature $\kappa$ [18] in Fig. 3, which we determined from 100 trajectories calculated with $N=50$ Euler steps with each trained model for each task, on samples from each respective training dataset, see Section S.X of the supplementary material. We plot the accumulated curvature $\kappa$, i.e., $\int_{0}^{\tau}\kappa(t)\mathrm{d}t$, to show both local changes in curvature and an informative overall curvature at $\tau=1$. We also plot the evaluation time-points $\mathbf{c}$ from the LRK solver learned for each task using (18). We can clearly see that the SE task has much lower curvature than all other tasks, reflecting the linear additive corruption model and explaining why even a single Euler step yields very good results here. The tasks then increase in total curvature in the following order: dereverberation, codec artifact removal, Mel vocoding and STFT phase retrieval, and BWE. Dereverberation and codec artifact removal have lower curvature and may have less inherent ambiguity than for instance BWE, and they profit more from using the signal-fidelity oriented PESQ-based loss (19) to train the LRK solver, see Tables IV(a) and IV(b).
For the LRK4 results on SE, shown in Tables II and III, we used a single set of LRK solver parameters trained on EWv2 with our main SFM model for SE. From these tables, we can see that these parameters generalize reasonably well to other models (16 ms SFM variants) and datasets (VB-DMD benchmark). Regarding cross-task evaluation, a complete evaluation, conducted by running inference for every task and model with every task-mismatched set of solver parameters, can be found in Section S.IX-I of the supplementary material. The results suggest that (1) the LRK solvers for SE, dereverberation, codec artifact removal and BWE transfer reasonably well to other tasks within this group, (2) the aforementioned solvers do not transfer well to PR and Mel vocoding, and (3) most interestingly, the solvers for PR and Mel vocoding conversely seem to be most universally applicable.
Another interesting question is how helpful LRK solvers are in settings with even lower NFE. To investigate this, we train one solver for each $N\in\{2,3,4,5\}$ for bandwidth extension, and compare each against the Euler solver at the same respective NFE. Note here that $N=1$ yields the plain Euler scheme with no learnable parameters. We show the results in Table VII, where we find that the biggest improvement over Euler occurs at $N=3$, though also at the cost of the largest decrease in PESQ.
Finally, regarding the time needed to training a LRK scheme, we find that this scales linearly with $N$, taking $2.3N$ hours for 25,000 on an NVIDIA RTX A6000. This yields around 11.5 hours for a 5-stage LRK solver, substantially less than the time taken to train the flow model DNN (62 hours) on the same GPU.
V-H Model weight compression for Mel vocoding
@@
V-I Model architecture ablations
[…]
FM -68.9 2.31 3.77 11.4%
Stream.FM 0.0 2.01 3.76 16.8%
+Bottleneck attn. +5.2NN 2.33 3.90 13.0%
+Causal win. bottl. attn. $\leq$ +5.2NN 2.04 3.79 16.6%
+Concat. fusion +62.9NN 2.00 3.76 17.5%
+Cum. GroupNorm +0.0 2.06 3.76 16.9%
[…]
As guidance for future work, we conduct ablation studies for some of our architectural choices in Section III-C, modifying only one component of the DNN at a time. We choose dereverberation as the task, motivated by the assumption that it especially profits from effective modeling of intricate long-range and phase relationships. We investigate the addition of non-causal bottleneck attention, the addition of causal windowed bottleneck attention with 64 past frames (1.024 seconds), using channel-concatenation instead of addition for skip connections, using a cumulative GroupNorm as in [48] instead of our SGBatchNorm, and using one instead of two residual blocks per layer. We use the Euler solver with 5 steps for a consistent comparison, and also list the non-causal FM model for comparison.
We show the resulting PESQ, DistillMOS, and WER values in Table VIII, along with the change in GFLOPs per second from each modification. We see that adding non-causal bottleneck attention strongly boosts the output quality, and is even slightly better in PESQ and DistillMOS than the non-causal FM model, which also uses bottleneck attention but on time-downsampled features. However, using causal windowed bottleneck attention instead to get a streaming-capable variant provides only a very minor quality increase. This supports our earlier claim that future information is very helpful in the dereverberation task, and justifies our choice of removing attention for simplicity and efficiency in our streaming-oriented setting. Using channel concatenation for skip connections does not yield any improvement while requiring $62.9N$ more GFLOPs per second, justifying our simpler choice of addition. Using a cumulative GroupNorm as in [48] instead of our SGBatchNorm slightly increases PESQ, but makes no clear overall positive or negative difference. We argue that in dynamic real-world acoustic environments (e.g., moving between rooms), a time-invariant norm such as ours is likely still preferable. Using one residual block instead of two clearly decreases all three metrics, supporting our use of two residual blocks.
@@ VI Conclusion
In this work, we have presented Stream.FM, a streaming generative method for general speech restoration tasks which can run in real-time on a consumer GPU. We detailed our architecture, inference scheme, and other optimizations needed to achieve this real-time capability, showed state-of-the-art performance for generative streaming methods through metrics and listening experiments, and proposed novel ideas to optimize quality or compute in a low-NFE setting. Our contributions and public codebase aim at closing the gap between generative and predictive speech restoration models in real-time settings, and at supporting the research community towards further developments in this field.
@@ VII Acknowledgements
Supplementary Material
S.VIII Model implementation and optimization
In this supplementary material, we list additional technical details for our Stream.FM models, which are not necessary for the method itself but are useful to implement real-time capable streaming inference.
S.VIII-A Functional frame-wise inference API
To implement our frame-wise inference scheme, we implement our model using PyTorch and PyTorch Lightning. We design an application programming interface (API) similar to the one proposed for CINs [16]. The official CINs implementation stores tracked states on each layer, making every layer stateful. This is problematic for our purposes, since we want to use only a single DNN instance $v_{\theta}$ with shared weights, but multiple state collections (one for each diffusion solver step).
@@
S.VIII-B Minimizing overhead
In initial experiments with our models for real-time streaming inference, we found that the processing time per frame quickly exceeded our 16 ms budget even for low NFE values such as 3. After careful inspection of profiler traces, we found that the processing was dominated by CPU-GPU overhead. To reduce this overhead as much as possible, we first tried using model compilation via torch.compile to merge CUDA kernels, which reduced the overhead to some extent but not enough to allow real-time inference. We then modified our code to use CUDA Graphs as implemented in PyTorch. By capturing the entire solver computation including the sequence of $N$ DNN calls in a single CUDA Graph and then replaying this CUDA Graph for every frame, we are able to realize up to 5 DNN calls per frame within the 16 ms time budget.
S.IX Further details on learned Runge-Kutta solvers
We provide further details on our proposed LRK solvers in the following.
S.IX-A Numerical properties of our learned RK solvers
Due to the constraints we impose on the learned RK coefficients, namely $\sum_{j}a_{ij}=c_{i},\sum_{i}b_{i}=1$ as described in the main paper, all our LRK methods are first-order consistent and have order of convergence of at least 1, by construction. Furthermore, as (consistent) Runge-Kutta schemes, they are (consistent) one-step methods and thus automatically have the property of zero-stability [32].
Through inspection of the learned coefficients, we find that higher convergence orders than 1 are not attained. This is expected, as such conditions require very specific algebraic relationships between the coefficients [15] that are extremely unlikely to be found by any learning-based method of determining the coefficients. This is, however, also acceptable since the solvers are intended to run at our fixed NFE budget and therefore at a fixed step $\Delta t$ (particularly in our case $\Delta t=1$), and are not intended to be general-purpose solvers.
We note here that, unlike the more general usage of RK ODE solvers which are called multiple times in sequence with some choice of step size $\Delta\tau$, our solvers are trained with $\Delta\tau=1$ for the fixed time horizon $\tau\in[0,1]$, and are intended to be called only once for each frame in order to guarantee a fixed NFE within our streaming constraints. Furthermore, the ODE induced by the learned flow matching velocity model $v_{\theta}$ is extremely high-dimensional and generally non-linear. This means that solver stability properties, which are based on the analysis of a low-dimensional linear test ODE and usually relate to long-term stability of repeated solver calls over unbounded time intervals [5], are less relevant for our single-call solvers than in a classical numerical setting.
Nonetheless, in line with classic literature on numerical ODE solvers [5], we can analyze and plot the region of absolute stability of each of our schemes, which can help characterize the solvers and may hold insights to be explored in future works. To this end, we introduce the stability function for Runge-Kutta schemes as in [5]:
$R(z):=1+z\mathbf{b}^{\top}(I-z\mathbf{A})^{-1}\mathbf{1}$ (21)
where $\mathbf{A},\mathbf{b}$ are taken from the coefficients of the LRK scheme and $\mathbf{1}$ indicates the vector $1,\ldots,1]$. The region of absolute stability is then the region of the complex plane where the condition $|R(z)|<1$ holds. We plot these regions in Fig. S5 for all LRK schemes learned with the SpeechBERTScore-based loss. The particularly large regions of absolute stability of the LRK solvers for the STFT phase retrieval and Mel vocoding tasks is interesting, and may relate to their high reliability across tasks which we find in Section S.IX-I. We reserve deeper investigations of possible relations for future work.
S.IX-B Learned coefficients for learned RK solvers
In this section, we list the parameters $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ of the learned Runge-Kutta ODE solver schemes, see Section III.E in the main paper, for each of the six tasks investigated in the main paper and for every solver learned with the SpeechBERTScore-based loss, Eq. (19) in the main manuscript. Note that values are rounded to three decimal points for clarity, and that the learned values at full precision always ensure the conditions we impose as described in the main paper (up to numerical floating-point error), e.g., $\sum_{i}b_{i}=1$.
S.IX-C Speech enhancement
@@
S.IX-E Codec post-filtering
@@
S.IX-G STFT phase retrieval